Find the equation of a straight line which oases through the point (2,-3) and is parallel to the line 2x + y = 6.


1
Expert's answer
2021-09-06T07:43:23-0400

The equation of the line "2x + y = 6" in slope-intercept form is


"y=-2x+6"

Hence


"slope=m=-2"

Then the equation of a line which is parallel to the line "y=-2x+6" is


"y=-2x+b"

If the line passes through the point "(2, -3)," then


"-3=-2(2)+b"

Solve for "b"


"b=1"

Then the equation of a line which is parallel to the line "2x + y = 6" and passes through the point "(2, -3)" is


"y=-2x+1"

Or


"2x+y=1"